OpenStudy (anonymous):

A rental car agency charges a flat fee of $110.00 plus $46.00 per day to rent a certain car. Another agency charges a fee of $70.00 plus $54.00 per day to rent the same car. Using a graphing calculator, find the number of days for which the costs are the same. Round your answer to the nearest whole day. 9 3 5 8

OpenStudy (anonymous):

@CalebBeavers

OpenStudy (anonymous):

Give me a second

OpenStudy (anonymous):

Do you have a graphing calculator? I dont, but i know how to solve it algebraically

OpenStudy (anonymous):

no i don't

OpenStudy (anonymous):

Alright well we'll do do it the algebraic way. the first agency's cost can be written as c=46d+110 an the second as c=54d+70. Do you see how i got that? c represents cost and d represents days

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Understand how i got those two?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Alright now we can solve the two linear equations, c=46d+110 c=54d+70 Move the variables to one side: c-46d=110 c-54d=70 Now how would you solve that?

OpenStudy (anonymous):

ok so I subtract both of them

OpenStudy (anonymous):

or add

OpenStudy (anonymous):

c-46d=110 c-54d=70 ------------ c-100=180

OpenStudy (anonymous):

this is what you're doing: adding these two together: c-46d=110 c-54d=70 but first you have to cancel out one of the variables so that you only have one to solve for. so how would we cancel out c?

OpenStudy (anonymous):

What would you mutiply the top equation by so that the c turns into -c?

OpenStudy (anonymous):

subtract c from both equations

OpenStudy (anonymous):

Focus on the top one for now, we need c to turn into -c so that when we add the c from the bottom one, it will equal 0 and we will just have d to solve for. so to do that we multiply the whole top equation by -1

OpenStudy (anonymous):

When we do that itll give us -c+46d=-110 c-54d=70 now when we add the equations together, the c will cancel out and well just have one variable that we can solve for

OpenStudy (anonymous):

So now add the equations together

OpenStudy (anonymous):

100 = -220

OpenStudy (anonymous):

Where did the d go? 46d+(-54d)=-8d and -110+70=-40 so were left with: -8d=-40

OpenStudy (anonymous):

Now solve for d and that is your answer

OpenStudy (anonymous):

oopps sorry forgot so then I will divide -8 from boths sides

OpenStudy (anonymous):

the answer would be 5

OpenStudy (anonymous):

Thats correct

OpenStudy (anonymous):

Thank you so much!!!! could you help me with another question?

OpenStudy (anonymous):

I could try, post it as a new one